# Shahriar Noroozizadeh

### v1

@@ C.4 Additional experiments on path-star geometry
Leaf–first-hop distance (Fig.˜16). In Fig.˜16 (an extended version of the heatmaps in Fig.˜6), entry $(i,j)$ is the cosine distance between the leaf embedding of path $i$ and the first-hop (indecipherable token) embedding of path $j$. Diagonal entries are low (a leaf lies close to its correct first hop), while off-diagonals are higher. This structure explains why the first-token-only objective succeeds in-weights (Fig.˜4-(Right)): the $k$-fold composition map reduces to a local geometric step in the learned representation.
Path–by–path distance (Fig.˜17). Instead of only analyzing the distance between the leaf and first nodes, we analyze the distance across all pairs of nodes in a path. For each pair of paths $(i,j)$, we compute the mean distance between all node embeddings on path $i$ and all node embeddings on path $j$.
@@
Figure 16: Evidence of global geometry in path-star task: Leaf-first-token cosine distance between node embeddings. We present again the heatmaps from Fig.˜6 with an additional heatmap in the middle. Entry $(i,j)$ is the mean cosine distance between the leaf token in (an unseen) path $i$ (row) and first/hardest token on (an unseen) path $j$ (col). Each heatmap corresponds to a different training objective: Left: trained on edges and path-finding task $(\mathcal{D}_{\texttt{edge}}\cup\mathcal{D}_{\texttt{path}}^{\to})$; Middle: trained on edges and hardest-token-finding task (not presented in the main paper); Right: edges only $(\mathcal{D}_{\texttt{edge}})$. We find that even on these unseen paths, the leaf and first token embeddings cluster together, regardless of whether path-finding supervision exists; however, the geometry is strongest with global supervision (however, in Mamba SSM, even the locally supervised model shows strong geometry; see Fig.˜9).
@@
Figure 17: Evidence of global geometry in path-star task: Pathwise average cosine{cosine} distance between node embeddings. While in Fig.˜16, we reported the cosine distance between the embeddings of the terminal nodes of a pair of (unseen) paths, here we consider an average over all nodes in those (unseen) paths. In particular, entry $(i,j)$ is the mean cosinecosine distance between nodes on path $i$ (row) and nodes on path $j$ (col). On the first two settings, as before, we find that the diagonal cells are low, implying closer embeddings. Thus, a geometry has emerged on unseen paths. However, unlike in Fig.˜6, we do not see any such signal when trained only on the edge memorization task.
UMAP projection (Fig.˜18, zoomed in version of Fig.˜6, middle). A different way to establish geometry is to directly visualize the paths. As done in Fig.˜6, we do this by projecting the token embeddings of all nodes with default UMAP settings (neighbors=15, min_dist=0.1). A zoomed in version of this is presented in Fig.˜18. We exclude the root token embedding in these projections since that is common to all paths. We find that different paths form well-separated clusters, and within each path cluster, nodes tend to arrange from leaf toward ${v}_{\tt{root}}$.
@@ Appendix D Edge supervision and training dynamics
There are tangential aspects of our training that are worth elaborating on: the role of reverse edges (§D.1), the role of pause tokens (§D.2), and the role of interleaving edge-memorization (§D.3).
@@ D.1.1 The critical role of reverse edges in the large path-star task
(Role of reverse edges) We find in Fig.˜19 that:
@@
    A Transformer trained on only the forward edges, struggles on both forward and reverse path-finding tasks (see the middle color in Fig.˜19).
@@
    A Transformer trained on only the reverse edges, achieves non-trivial accuracy on the reverse path-finding task; however, it fails on the forward path-finding task (see the third color in Fig.˜19).
@@
We note that the asymmetry between forward and reversed path-generation tasks stems from their algorithmic complexity. The reversed path generation is algorithmically trivial on path-star graphs because each node has a unique predecessor along any target path—the model simply needs to follow the unique backward edges. Forward generation, however, requires planning (examine each outgoing path) or lookahead (track the reverse path without explicit chain-of-thought and reverse it). Indeed, for the in-context task of B&N’24, the model fails on the forward task, but strikingly succeeds on the reverse task, as corroborated in Fig.˜20 (right). Even in the in-weights setting Fig.˜20 (left), the reverse path is generally quicker to learn and yields higher accuracy.
Figure 19: Mixed edge supervision enables forward path generation while forward-only fails due to reversal curse. Exact-match accuracy on held-out leaves for multiple path-star graphs (varying degree $d$ and path length $\ell$). As established, training on mixed edges $\mathcal{D}_{\texttt{edge}}$ yields high non-trivial forward accuracy across graphs. But training on forward-only $\mathcal{D}_{\texttt{edge}}^{\to}$ fails on both the forward and reverse tasks. This is indicative of the reversal curse. With backward-only edges ($\mathcal{D}_{\texttt{edge}}^{\leftarrow}$) the model attains high accuracy primarily on reverse path generation for smaller graphs, This can be reconciled by noting that generating the reverse path is an easier retrieval task. Random forward path accuracy is $1/d$.
@@ D.1.2 Tiny graphs with uni-directional edge-memorization
In Fig.˜21, we revisit our tiny graphs in §C.3, and examine the embeddings of a Transformer when it memorizes only one direction of the edges. We find that a geometry still arises, although a bit weaker.
Figure 21: Embeddings of a Transformer with bi-directional vs. uni-direction edge memorization. With our smaller graphs, we find a geometry arise regardless of whether the model is made to memorize both or only one direction of each edge. However, the geometry is weaker (e.g., for the grid graph) under uni-directional memorization.
@@ D.2 Pause tokens for computational slack
Fig.˜22 shows that adding a short sequence of pause tokens after the prompt reliably boosts exact-match accuracy across graphs, for a given amount of training time. Increasing the number of pause tokens increases speed of convergence.
@@
However, we confirm that even in this regime, our models do achieve a high peak accuracy (see Fig.˜23). The fact that the composition task is learnable in this regime implies that the edge-pretrained model must have come with an adequate global geometry despite being trained only on local supervision (Section˜3.2).
Figure 23: Locally supervised model succeeds at path-finding task. We report the peak accuracy under finetuning an edge-memorizing model on our path-finding task of §2. We emphasize that this accuracy value is only reached momentarily, and typically deteriorates quickly during further finetuning. Nevertheless, this suggests that local supervision alone was adequate to synthesize a global geometry (Section˜3.2). Learning rates of the two phases are given in §B.3.
D.4 Learning order of tokens
For clarity, in Fig.˜24, we provide a side-by-side contrast between the order in which tokens are learned in the in-weights setup (with next-token prediction) and the in-context setup (with multi-token prediction).
@@ E.1 Empirical failure of composition learning under associative memory
E.2 Succinctness does not break the tie: Proof of Proposition 1
@@
Proof of Main Proposition˜1
@@
This is straightforward to see for $\ell_{2}$ norm based complexity based on Proposition˜3. For the bit complexity estimates, from Proposition˜2, we know that associative memory costs $|{V}|\log|E|$ many bits; for the geometry, we need to pin down the values of the embedding dimension $m$ and the cell count $\Delta$.
The path-star graph can be embedded such that each path is stored along a unique dimension (thus totally $m=d$ dimensions), and each dimension can be gridded into $\ell$ many cells. This requires a bit complexity of $|{V}|d\log\ell$. For a more ambitious geometry, we could be further squeeze this into $\log d$ dimensions while still keeping the paths well-separated (by the Johnson–Lindenstrauss lemma), resulting in $|{V}|\log d\log\ell$ bits. This is still greater than the cost of associative memory which is approximately $|{V}|(\log d\ell)=|{V}|(\log d+\log\ell)$.
@@
E.3 Node2Vec can represent a form of associative memory
In the standard associative memory view that we have discussed, associations are represented through the function ${\bf{\Phi}}(v)^{T}{\bf{W}}_{\tt assoc}{\bf{\Phi}}(u)$. However, dual encoder models like Node2Vec can only represent functions of the form ${\bf{\Phi}}(v)^{T}{\bf{\Phi}}(u)$. While this precludes the form of associative memory we care about, it still allows a contrived form of associative memory when there is a sufficiently large embedding dimensionality. Below we show that given a set of edges $E$, and a dimensionality of $|E|$, we can construct embeddings such that the dot product of two adjacent vertices are high, but any non-adjacent vertices have zero dot product. [^6] We suspect this should also be a lower bound i.e., such a large dimensionality must be needed to represent associatively in Node2Vec. In this sense, only local information is captured, whereas no global geometry is.
@@ E.3 Node2Vec can represent a form of associative memory
Dual encoder models like Node2Vec can represent memory associatively with an embedding dimensionality of $|E|$ where $E$ is the set of pairwise associations.
@@
Assume that each node is embedded in an $|E|$-dimensional space, where the $i$th dimension corresponds to the $i$th edge. Then, for edge $(u,v)$ we assume that the embedding of $u$ and $v$ both are set to $1$ along the dimension corresponding to $(u,v)$. Notationally, if the edges are indexed as $1,2,\ldots$, then ${\bf{\Phi}}(u)_{i}=\mathbf{1}[u\in e_{i}]$, where $\bf{1}$ is the indicator function. Then, we have that ${\bf{\Phi}}(u)\cdot{\bf{\Phi}}(v)=\mathbf{1}[(u,v)\in E]$. Thus, the dot products capture only local information. Any two non-adjacent vertices have a zero dot-product. No worthwhile geometry exists. ∎
@@ F.1 Challenges of analyzing the dynamics
Unlike previously-studied dynamics which simplify nicely, this system may behave in one of many ways. For one, it may simply diverge, but if we are a bit lucky, it may at least converge in direction (like in logistic regression [146]); but then, this direction then may be degenerate—an all-one representation could potentially be a stable direction—and perhaps nice directions are visible only if we analyze with early-stopping. One way to get a handle of this would have been to show that in the limit, we have ${\bf{C}}(t)\to 0$; solving this could then spell out the (limit) probability matrix $\bf{P}$}, if not the inner products ${\bf{V}}{\bf{V}}^{T}$ themselves. However, we find that $\bf{C}$} cannot be zero as that would require the self-probability term $p(i,i)=0$, which is infeasible. This closes all obvious analytical routes to understanding this system, so we turn to an empirical study.

### v2

@@ C.4 Additional experiments on path-star geometry
Leaf–first-hop distance (Fig.˜25). In Fig.˜25 (an extended version of the heatmaps in Fig.˜6), entry $(i,j)$ is the cosine distance between the leaf embedding of path $i$ and the first-hop (indecipherable token) embedding of path $j$. Diagonal entries are low (a leaf lies close to its correct first hop), while off-diagonals are higher. This structure explains why the first-token-only objective succeeds in-weights (Fig.˜4-(Right)): the $k$-fold composition map reduces to a local geometric step in the learned representation.
Path–by–path distance (Fig.˜26). Instead of only analyzing the distance between the leaf and first nodes, we analyze the distance across all pairs of nodes in a path. For each pair of paths $(i,j)$, we compute the mean distance between all node embeddings on path $i$ and all node embeddings on path $j$.
@@
Figure 25: Evidence of global geometry in path-star task: Leaf-first-token cosine distance between node embeddings. We present again the heatmaps from Fig.˜6 with an additional heatmap in the middle. Entry $(i,j)$ is the mean cosine distance between the leaf token in (an unseen) path $i$ (row) and first/hardest token on (an unseen) path $j$ (col). Each heatmap corresponds to a different training objective: Left: trained on edges and path-finding task $(\mathcal{D}_{\texttt{edge}}\cup\mathcal{D}_{\texttt{path}}^{\to})$; Middle: trained on edges and hardest-token-finding task (not presented in the main paper); Right: edges only $(\mathcal{D}_{\texttt{edge}})$. We find that even on these unseen paths, the leaf and first token embeddings cluster together, regardless of whether path-finding supervision exists; however, the geometry is strongest with global supervision (however, in Mamba SSM, even the locally supervised model shows strong geometry; see Fig.˜15).
@@
Figure 26: Evidence of global geometry in path-star task: Pathwise average cosine{cosine} distance between node embeddings. While in Fig.˜25, we reported the cosine distance between the embeddings of the terminal nodes of a pair of (unseen) paths, here we consider an average over all nodes in those (unseen) paths. In particular, entry $(i,j)$ is the mean cosinecosine distance between nodes on path $i$ (row) and nodes on path $j$ (col). On the first two settings, as before, we find that the diagonal cells are low, implying closer embeddings. Thus, a geometry has emerged on unseen paths. However, unlike in Fig.˜6, we do not see any such signal when trained only on the edge memorization task.
UMAP projection (Fig.˜27, zoomed in version of Fig.˜6, middle). A different way to establish geometry is to directly visualize the paths. As done in Fig.˜6, we do this by projecting the token embeddings of all nodes with default UMAP settings (neighbors=15, min_dist=0.1). A zoomed in version of this is presented in Fig.˜27. We exclude the root token embedding in these projections since that is common to all paths. We find that different paths form well-separated clusters, and within each path cluster, nodes tend to arrange from leaf toward ${v}_{\tt{root}}$.
@@ Appendix D Edge supervision and training dynamics
There are tangential aspects of our training that are worth elaborating on: the roles of reverse edges (§D.1), of pause tokens (§D.2), of interleaving edge-memorization (§D.3), of weight-tying (§D.4), and of the learning order of tokens (§D.5).
@@ D.1.1 The critical role of reverse edges in the large path-star task
(Role of reverse edges) We find in Fig.˜28 that:
@@
    A Transformer trained on only the forward edges, struggles on both forward and reverse path-finding tasks (see the middle color in Fig.˜28).
@@
    A Transformer trained on only the reverse edges, achieves non-trivial accuracy on the reverse path-finding task; however, it fails on the forward path-finding task (see the third color in Fig.˜28).
@@
We note that the asymmetry between forward and reversed path-generation tasks stems from their algorithmic complexity. The reversed path generation is algorithmically trivial on path-star graphs because each node has a unique predecessor along any target path—the model simply needs to follow the unique backward edges. Forward generation, however, requires planning (examine each outgoing path) or lookahead (track the reverse path without explicit chain-of-thought and reverse it). Indeed, for the in-context task of B&N’24, the model fails on the forward task, but strikingly succeeds on the reverse task, as corroborated in Fig.˜29 (right). Even in the in-weights setting Fig.˜29 (left), the reverse path is generally quicker to learn and yields higher accuracy.
Figure 28: Mixed edge supervision enables forward path generation while forward-only fails due to reversal curse. Exact-match accuracy on held-out leaves for multiple path-star graphs (varying degree $d$ and path length $\ell$). As established, training on mixed edges $\mathcal{D}_{\texttt{edge}}$ yields high non-trivial forward accuracy across graphs. But training on forward-only $\mathcal{D}_{\texttt{edge}}^{\to}$ fails on both the forward and reverse tasks. This is indicative of the reversal curse. With backward-only edges ($\mathcal{D}_{\texttt{edge}}^{\leftarrow}$) the model attains high accuracy primarily on reverse path generation for smaller graphs. This can be reconciled by noting that generating the reverse path is an easier retrieval task. Random forward path accuracy is $1/d$.
@@ D.1.2 Tiny graphs with uni-directional edge-memorization
In Fig.˜30, we revisit our tiny graphs in §C.3, and examine the embeddings of a Transformer when it memorizes only one direction of the edges. We find that a geometry still arises, although a bit weaker.
Figure 30: Embeddings of a Transformer with bi-directional vs. uni-directional edge memorization. With our smaller graphs, we find a geometry arise regardless of whether the model is made to memorize both or only one direction of each edge. However, the geometry is weaker (e.g., for the grid graph) under uni-directional memorization.
@@ D.2 Pause tokens for computational slack
Fig.˜31 shows that adding a short sequence of pause tokens after the prompt reliably boosts exact-match accuracy across graphs, for a given amount of training time. Increasing the number of pause tokens increases speed of convergence.
@@
However, we confirm that even in this regime, our models do achieve a high peak accuracy (see Fig.˜32). The fact that the composition task is learnable in this regime implies that the edge-pretrained model must have come with an adequate global geometry despite being trained only on local supervision (Section˜3.1).
Figure 32: Locally supervised model succeeds at path-finding task. We report the peak accuracy under finetuning an edge-memorizing model on our path-finding task of §2. We emphasize that this accuracy value is only reached momentarily, and typically deteriorates quickly during further finetuning. Nevertheless, this suggests that local supervision alone was adequate to synthesize a global geometry (Section˜3.1). Learning rates of the two phases are given in §B.3.
D.4 Effect of weight-untying
As discussed in §4.4, weight-untied models do not produce an apparent geometry or exhibit high node-to-node cosine-similarities, since it corresponds to SVD directions. We demonstrate this for the Transformers in Fig.˜11 and Fig.˜34 and Node2Vec in Fig.˜10 and Fig.˜33.
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 33: Node-node cosine similarity vs. adjacency matrix for weight-untied Node2Vec. As discussed in §4.4, observe that the node-to-node cosine-similarity matrix appears to be a bland near-zero matrix, lacking any information about the multi-hop connectivity of the underlying graph.
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 34: Node-node cosine similarity vs. adjacency matrix for weight-untied Transformer. As discussed in §4.4, observe that the node-to-node cosine-similarity matrix appears to be a bland near-zero matrix, lacking any information about the multi-hop connectivity of the underlying graph.
D.5 Learning order of tokens
For clarity, in Fig.˜35, we provide a side-by-side contrast between the order in which tokens are learned in the in-weights setup (with next-token prediction) and the in-context setup (with multi-token prediction).
@@ E.1 Empirical failure of composition learning under associative memory
E.2 Succinctness does not break the tie: Proof of Proposition 3.2
@@
Proof of Main Section˜3.2
@@
This is straightforward to see for $\ell_{2}$ norm based complexity based on Proposition˜2. For the bit complexity estimates, from Proposition˜1, we know that associative memory costs $|{V}|\log|E|$ many bits; for the geometry, we need to pin down the values of the embedding dimension $m$ and the cell count $\Delta$.
The path-star graph can be embedded such that each path is stored along a unique dimension (thus totally $m=d$ dimensions), and each dimension can be gridded into $\ell$ many cells. This requires a bit complexity of $|{V}|d\log\ell$. For a more ambitious geometry, we could further squeeze this into $\log d$ dimensions while still keeping the paths well-separated (by the Johnson–Lindenstrauss lemma), resulting in $|{V}|\log d\log\ell$ bits. This is still greater than the cost of associative memory which is approximately $|{V}|(\log d\ell)=|{V}|(\log d+\log\ell)$.
@@
E.3 Weight-tied dual-encoder models can represent certain special or contrived forms of associative memory
In the standard associative memory view that we have discussed, associations are represented through the function ${\bf{\Phi}}(v)^{T}{\bf{W}}_{\tt assoc}{\bf{\Phi}}(u)$. However, weight-tied dual encoder models like Node2Vec can only represent functions of the form ${\bf{\Phi}}(v)^{T}{\bf{\Phi}}(u)$. When weight-tied, this precludes the form of associative memory we care about; but it still allows two forms of associative memories under special conditions: if the graph is bipartite (Proposition˜3) or if the embedding dimensionality is sufficiently large (in which case, the construction is contrived; see Proposition˜5). We can also show that non-weight-tied models can represent any graph associatively (see Proposition˜4). In all these results below, we will show that our constructions achieve a logit of $1$ on adjacent vertices, but a logit of 0 on any non-adjacent vertices. In this sense, only local associations are captured, whereas no global geometry is.
@@
Weight-tied dual encoder models of the form ${\bf{\Phi}}(v)\cdot{\bf{\Phi}}(u)$ can store a graph $\mathcal{G}$ associatively when the graph is bipartite.
@@
Assume that the graph partition consists of vertices ${V}_{1}$ and ${V}_{2}$. The construction is to represent one set of vertices with a one-hot representation, and the other by a one-hot adjacency vector. That is, consider an embedding ${\bf{\Phi}}:{V}\to\mathbb{R}^{|{V}_{1}|}$ that has dimensionality $|{V}_{1}|$. For any $v_{i}\in{V}_{1}$, assume ${\bf{\Phi}}(v)$ is such that it is $1$ only at index $i$. For $w\in{V}_{2}$, ${\bf{\Phi}}(v)$ at index $i$ is $1$ if and only if $v$ is adjacent to $v_{i}$ i.e., ${\bf{\Phi}}(v)_{i}=\mathbf{1}[(w,v_{i})\in E]$.
Thus, if $w$ and $v$ belong in the same partition ${\bf{\Phi}}(v)\cdot{\bf{\Phi}}(w)=0$; if in different partitions, ${\bf{\Phi}}(v)\cdot{\bf{\Phi}}(w)=\mathbf{1}[(w,v)\in E]$. ∎
Proposition 4.
Weight-untied dual encoder models of the form ${\bf{\Phi}}_{1}(v)\cdot{\bf{\Phi}}_{2}(u)$ can store any graph associatively when the embedding dimension is at least as large as the vertex count. Furthermore, this can be approximately learned in one-step of gradient descent under the correlation loss $\sum_{(u,v)\in E}({\bf{\Phi}}_{1}(u))\cdot{\bf{\Phi}}_{2}(v)$ with one tower frozen.
Proof.
The proof parallels that of Proposition˜3. We set one of the embeddings to a one-hot representation of the vectors; the other is set to embed the adjacency row. In matrix notation, the resulting model (which outputs a vector of logits for a given input $\bf{u}$}) looks like $f({\bf{u}})={\bf{W}}_{2}{\bf{W}}_{1}{\bf{u}}={\bf{A}}{\bf{I}}{\bf{u}}={\bf{A}}{\bf{u}}$.
To show that this can be learned in one-step of gradient descent, consider ${\bf{\Phi}}_{1}$ initialized randomly. Given sufficiently high embedding size, these embeddings are orthogonal. Under the correlation loss, if one were to freeze ${\bf{\Phi}}_{1}$, ${\bf{\Phi}}_{2}(v)$ will be updated in the direction of $\sum_{u\in\texttt{nbr}(v)}{\bf{\Phi}}_{2}(u)$, which is equivalent to the adjacency row up to a rotation. For a sufficiently large learning rate, the resulting model would effectively store $\bf{A}$} up to a rotation defined by ${\bf{\Phi}}_{1}$. ∎
Now, for a generic graph, and for a weight-tied dual encoder model, we show that Node2Vec can represent a contrived form of associative memory when there is a much larger embedding dimensionality scaling with edge count.  [^14] We suspect this should also be a lower bound i.e., such a large dimensionality must be needed to represent associatively in Node2Vec.
Proposition 5.
Weight-tied dual encoder models of the form ${\bf{\Phi}}(v)\cdot{\bf{\Phi}}(u)$ can store a graph associatively given an embedding dimensionality of $|E|$ where $E$ is the set of pairwise associations.
Proof.
For each node, we take its corresponding row from the “incidence” matrix. In other words, the embedding is $|E|$-dimensional, where the $i$th dimension corresponds to the $i$th edge; for edge $(u,v)$, the embedding of $u$ and $v$ both are set to $1$ along the dimension corresponding to $(u,v)$. Notationally, if the edges are indexed as $1,2,\ldots$, then ${\bf{\Phi}}(u)_{i}=\mathbf{1}[u\in e_{i}]$, where $\bf{1}$ is the indicator function. Then, we have that ${\bf{\Phi}}(u)\cdot{\bf{\Phi}}(v)=\mathbf{1}[(u,v)\in E]$. Thus, the dot products capture only local information. ∎
@@ F.1 Challenges of analyzing the dynamics
Unlike previously-studied dynamics which simplify nicely, this system may behave in one of many ways. For one, it may simply diverge, but if we are a bit lucky, it may at least converge in direction (like in logistic regression [162]); but then, this direction then may be degenerate—an all-one representation could potentially be a stable direction—and perhaps nice directions are visible only if we analyze with early-stopping. One way to get a handle on this would have been to show that, in the limit, we have ${\bf{C}}(t)\to 0$; solving this could then spell out the (limit) probability matrix $\bf{P}$}, if not the inner products ${\bf{V}}{\bf{V}}^{T}$ themselves. However, we find that $\bf{C}$} cannot be zero as that would require the self-probability term $p(i,i)=0$, which is infeasible. This closes all obvious analytical routes to understanding this system, so we turn to an empirical study.